OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system? Part 2: Solve the following system and show all of your work. 3y - 2x = 11 y + 2x = 9

OpenStudy (texaschic101):

that means there is infinite number of solutions. If it came out like 2 = 4, then there would be no solutions. y + 2x = 9 y = -2x + 9 now sub -2x + 9 in for y in the other equation 3y - 2x = 11 3(-2x + 9) - 2x = 11 -- distribute through the parenthesis -6x + 27 - 2x = 11 -- subtract 27 from both sides -6x - 2x = 11 - 27 -- combine like terms -8x = - 16 -- divide by -8 x = -16/-8 x = 2 now sub 2 in for x in either of the equations to find y y + 2x = 9 y + 2(2) = 9 y + 4 = 9 y = 9 - 4 y = 5